Grade 11Mechanics

A ball is thrown vertically upwards from a 85m high building and it reaches the ground after 5s.calculate initial and final velocities of the body

Divyanshu yadav
9 Years agoGrade 11

1 Answer

Agastya Varma
9 Years ago
In this question take downwards as positive.
Sy = uyt+ 1/2ayt2   ( here we are divinding the values into their y components)
85= u(5) + 1/2(10)(52)
U= -8m/s     – initial velocity.      ( we got -8 because we took downwards as positive)
 
V= u + at
   = -8 + (10)(5)  (we are taking -8 because we took downwards as positive.
   = 42m/s